The range of projectile is 4 times of the height of projectile at an angle of?



A. 75



B. 76



C. 60



D. 90

1
Expert's answer
2022-03-16T09:41:36-0400
"R=4H_{\max}"

"\frac{2v_0^2\sin\theta\cos\theta}{g}=4*\frac{v_0^2\sin^2\theta}{2g}"

"\tan\theta=1"

"\theta=45^\circ"